# Learning to Pursue AC Optimal Power Flow Solutions with Feasibility Guarantees

### v1

@@ IV-A Dataset Generation and Training
Figure 3: Overvoltage events and number of nodes impacted with the considered simulation setup, when no control actions are implemented.
[…]
@@
(b) Online NN-SGF in Algorithm 2 and Fig. 1(left).
(c) Neural network trained to emulate $\textsf{U}^{\textsf{lm}}(\boldsymbol{s}_{l},\boldsymbol{\theta})$.
Figure 4: Highest voltage profile and number of nodes experiencing overvoltages with different optimization methods : (a) Batch optimization (BO) using IPOPT; (b) the proposed NN-SGF (8), implemented in an online feedback configuration as in Figure 1(left); (c) method where a neural network is trained to emulate BO solutions (NN-BO).
@@
NN-SGF (s2) 60,000 1.7 ×$\times 10^{-6}$
NN-BO (s3) 80,000 8.3 ×$\times 10^{-5}$
IV-B Voltage regulation and over-voltage duration
Figures 4 and 5 illustrate how different strategies perform in regulating voltage magnitudes within the bounds $[0.95,1.05]$ p.u.; Figure 4 considers the maximum voltage profile across the system at every time step, as well as the number of nodes that experience overvoltages. The proposed NN–SGF method maintains voltages tightly bounded across all monitored nodes, with only brief and mild excursions slightly above 1.05 p.u. These short-duration deviations are well within the tolerance accepted by distribution utilities and do not compromise protection schemes. The BO method, which solves the full AC OPF offline to convergence, is used as a benchmark. The NN-BO approach, trained to emulate the BO setpoints directly, exhibits significantly more overvoltage excursions than the NN-SGF, reflected in its higher $\max T_{1.05}$ and $\mathrm{mean}\ T_{1.05}$ as shown in Figure 5. This confirms that approaches that attempts to learn solutions to the OPF directly cannot ensure feasibility. Importantly, NN-SGF delivers effective online voltage regulation without any iterative optimization and even outperforms widely used schemes such as Volt/Var Control (VVC) and online primal dual methods investigated in [24], which typically suffer from slower response or larger transient violations. Overall, NN-SGF strikes the best balance among voltage compliance, computational efficiency, and system protection, all without introducing operational concerns.
Figure 5: Overvoltage duration for the three strategies (s1)(s3) compared in Figure (4). Additionally, we consider the no control (NC) setup as in Figure 3 and the SFG.
IV-C Computational times
[…]
@@ IV-C Computational times
[…]
@@
Method Online step Offline implementation Offline
@@
We also consider the offline implementation. Here, the “Offline solution” refers to Figure 1(center), where each iteration involves one evaluation of the SGF (resp., the NN-SGF) and one solution to the PF. Again, the PF equations are solved using pandapower, and the average execution time of pandapower was 0.021 seconds. On average, the proposed scheme implemented in an offline fashion required less than 10 iterations, yielding the upper bounds provided in Table II.
We note that the proposed NN-SGF requires measurements of the voltages, while the BO and NN-BO require measurements of all the loads in the network; therefore, the actual time required by BO and NN-BO is much larger in practice [36].
@@ V Theoretical Analysis
Assumption 2 (Regularity of isolated solutions).
Assume that (II-B) is feasible and let $\boldsymbol{u}^{*}$ be a local minimizer and an isolated KKT point for (II-B), for given $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$. Assume that:
@@
ii) The maps $\boldsymbol{u}\mapsto C_{p}(\boldsymbol{u})$, $\boldsymbol{u}\mapsto C_{v}(|\boldsymbol{v}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})|)$, $\boldsymbol{u}\mapsto|\boldsymbol{v}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})|$, and $\boldsymbol{u}\mapsto|\boldsymbol{i}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})|$ are twice continuously differentiable over $\mathcal{B}(\boldsymbol{u}^{*},r_{1})$, and their Hessian matrices are positive semi-definite at $\boldsymbol{u}^{*}$.
iii) The Hessian $\nabla^{2}C_{p}(\boldsymbol{u}^{*})$ is positive definite. $\Box$
@@
Assumption 3 (Jacobian errors).
$\exists~{}E_{v}<+\infty,E_{J_{v}}<+\infty$ such that $\||\boldsymbol{v}(\boldsymbol{u};\boldsymbol{s}_{l})|-(\boldsymbol{\Gamma}_{v}\boldsymbol{u}+\bar{\boldsymbol{v}}(\boldsymbol{s}_{l}))\|\leq E_{v}$ and $\|\boldsymbol{\Gamma}_{v}-J_{v}(\boldsymbol{u},\boldsymbol{s}_{l})\|\leq E_{J_{v}}$ for any $\boldsymbol{u}\in\mathcal{B}(\boldsymbol{u}^{*},r_{1})$. $\Box$
@@
$\exists~{}\epsilon_{n}<+\infty$ such that $\|\boldsymbol{n}\|\leq\epsilon_{n}$. Let $\epsilon_{v}<+\infty$ and ${}\epsilon_{i}<+\infty$ such that $\|\boldsymbol{n}_{v}\|\leq\epsilon_{v}$ and $\|\boldsymbol{n}_{i}\|\leq\epsilon_{i}$, respectively. $\Box$
@@
Since the line currents $\boldsymbol{i}$ can be computed from $\boldsymbol{v}$ via Ohm’s Law, Assumption 3 implies that $\exists~{}E_{i}<+\infty,E_{J_{i}}<+\infty$ such that $\||\boldsymbol{i}(\boldsymbol{u};\boldsymbol{s}_{l})|-(\boldsymbol{\Gamma}_{i}\boldsymbol{u}+\bar{\boldsymbol{i}}(\boldsymbol{s}_{l}))\|\leq E_{i}$ and $\|\boldsymbol{\Gamma}_{i}-J_{i}(\boldsymbol{u},\boldsymbol{s}_{l})\|\leq E_{J_{i}}$ for any $\boldsymbol{u}\in\mathcal{B}(\boldsymbol{u}^{*},r_{1})$. Assumptions 3-4 are motivated by the fact that the error of the linear approximation is small in a neighborhood of the optimizer [27, 24]), and that in realistic monitoring and SCADA systems, the measurement of the voltage magnitudes are affected by a negligible error [36]. Lastly, Assumption 5 follows from the approximation capabilities of neural networks over compact sets [40, 41].
@@
$:=\arg\min_{\boldsymbol{z}\in\mathbb{R}^{2G}}\|\boldsymbol{z}+\nabla C_{p}(\boldsymbol{u})+(\boldsymbol{J}_{v}(\boldsymbol{u};\boldsymbol{s}_{l})+\boldsymbol{\Psi}_{v})^{\top}\nabla C_{v}(\boldsymbol{\nu})\|^{2}$
[…]
@@
Assumption 6 (Regularity).
For any $\boldsymbol{u}\in\mathcal{B}(\boldsymbol{u}^{*},r_{1})$, and any $\boldsymbol{\Psi}_{v}$, $\boldsymbol{\Psi}_{i}$, and $\mathbf{e}$ satisfying Assumptions 3-4, the problem (15) is feasible, and satisfies the Mangasarian-Fromovitz Constraint Qualification and the constant-rank condition [35]. $\Box$
@@
Lemma V.1 (Lipschitz continuity).
Let Assumption 6 hold, and assume that $\boldsymbol{u}\mapsto C_{p}(\boldsymbol{u})$, $\boldsymbol{\nu}\mapsto C_{v}(\boldsymbol{\nu})$ are twice continuously differentiable over $\mathcal{B}(\boldsymbol{u}^{*},r_{1})$ and for any $\boldsymbol{\nu}$ Then:
@@
Lemma V.1 follows from [35, Theorem 3.6], and by the compactness of the sets $\mathcal{E}_{n}$, $\mathcal{E}_{v}$ and $\mathcal{E}_{i}$. To state the main result, recall that $\boldsymbol{u}^{*}$ is a local optimizer of (II-B). Define $\boldsymbol{\xi}^{*}:=H(\boldsymbol{u}^{*};\boldsymbol{s}_{l})$, $\boldsymbol{J}_{F}:=\frac{\partial F(\boldsymbol{u},H(\boldsymbol{u};\boldsymbol{s}_{l}),\boldsymbol{\theta})}{\partial\boldsymbol{u}}\mid_{\boldsymbol{u}=\boldsymbol{u}^{*}}$, $e_{1}:=-\lambda_{\max}(\boldsymbol{J}_{F})$, and $e_{2}:=-\lambda_{\min}(\boldsymbol{J}_{F})$. Then, from [23], we can write the dynamics as $F(\boldsymbol{u},H(\boldsymbol{u};\boldsymbol{s}_{l}),\boldsymbol{\theta})=\boldsymbol{J}_{F}(\boldsymbol{u}-\boldsymbol{u}^{*})+g(\boldsymbol{u})$, where $g$ satisfies $\|g(\boldsymbol{u})\|\leq L\|\boldsymbol{u}-\boldsymbol{u}^{*}\|^{2}$, $\forall\boldsymbol{u}\in\mathcal{B}(\boldsymbol{u}^{*},r_{2})$, for some $L>0$ and $r_{2}>0$. Define $r:=\min\{r_{1},r_{2}\}$ and $s_{\min}$ as: $s_{\min}=0$ if $r\geq\frac{e_{1}}{L}$, and $s_{\min}=1-\frac{rL}{e_{1}}$ if $r\geq\frac{e_{1}}{L}$. We are ready to state the main result.
Theorem V.2 (Stability and convergence).
Consider the OPF problem (II-B) satisfying Assumption 1. Let Assumptions 35 hold for the linear model and the training, and let Assumption 6 hold for (11). Let $\boldsymbol{u}(t)$, $t\geq t_{0}$, be the unique trajectory of (8). Let $\epsilon:=\ell_{J_{v}}E_{J_{v}}+\ell_{J_{i}}E_{J_{i}}+\ell_{n}\epsilon_{n}+\epsilon^{\textsf{NN}}$, and assume that the set :=$\mathcal{R}:=\left\{s:s_{\min}<s\leq 1,~{}e_{1}^{-3}e_{2}LE<s-s^{2}\right\}$ is not empty. Then, for any $s\in\mathcal{R}$, it holds that
@@
Proposition V.3 (Practical forward invariance).
@@
$\mathcal{S}_{e,v}$ :=$:=\{\boldsymbol{u}\in\mathcal{C}:\underline{V}_{e}\leq|\boldsymbol{v}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{V}_{e}\}$
$\mathcal{S}_{e,i}$ :=$:=\{\boldsymbol{u}\in\mathcal{C}:|\boldsymbol{i}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{I}_{e}\}$

### v2

@@ IV-A Dataset Generation and Training
Figure 3: Overvoltage events under the considered simulation setup when no control actions are implemented. Evaluated over the monitored buses $\mathcal{M}=\{93,54,85,84,48,39,91,71,62,67\}$.
For the method (s3) used for comparison, the training of an FNN $\mathcal{F}_{\textsf{batch}}^{\textsf{NN}}:(\boldsymbol{s}_{l},\boldsymbol{\theta})\mapsto\boldsymbol{u}^{\ast}$ to emulate solutions to the BO was computationally heavier as we had to increase the size of the training set to obtain acceptable performance. In fact, by using 60,000 training points (as for the NN-SGF), the performance of the NN-BO was poor; we then resorted to 20,000 additional training points for the NN–BO to obtain a smaller test error. This increased number of data points reflects the higher sample complexity of directly learning the nonconvex AC–OPF solution map. The training set consists of inputs $\boldsymbol{s}_{l}^{(k)},\boldsymbol{\theta}^{(k)}$, with corresponding outputs $\boldsymbol{u}^{\ast(k)}$ obtained by solving the AC OPF using IPOPT. To generate the dataset, we sample $N_{\text{cases}}=8{,}000$ time instants $t_{n}\sim\mathcal{U}(\text{06:00},\text{20:00})$, record the uncontrollable loads $\boldsymbol{s}_{l}(t_{n}):=(\boldsymbol{p}_{l}(t_{n}),\boldsymbol{q}_{l}(t_{n}))$ and inverter limits $\boldsymbol{p}_{\max}(t_{n})$, and apply perturbations $\boldsymbol{s}_{l}^{(n,k)}:=(1+\epsilon_{k})\boldsymbol{s}_{l}(t_{n})$ with $\epsilon_{k}\sim\mathcal{U}(-0.05,0.05)$ for $k=1,\dots,10$, introducing up to $\pm 5\%$ variability to enable localized sampling of the solution space without violating OPF feasibility at $t_{n}$. Solving the AC OPF for each perturbed point yields the target $\boldsymbol{u}^{\ast(n,k)}\in\textsf{U}^{\textsf{lm}}(\boldsymbol{s}_{l}^{(n,k)},\boldsymbol{\theta}))$, giving $N_{\text{train}}=80{,}000$ training pairs. The loss minimized is $\mathcal{L}(\theta)=\frac{1}{N_{\text{train}}}\sum_{i=1}^{N_{\text{train}}}\left\|\mathcal{F}_{\textsf{batch}}^{\textsf{NN}}(\boldsymbol{s}_{l}^{(i)},\boldsymbol{\theta};\theta)-\boldsymbol{u}^{\ast(i)}\right\|_{2}^{2}$. Using the same laptop with a 16-core Intel Core Ultra 7 255H CPU and 32 GB of RAM, training NN–BO takes approximately $110$ minutes.
@@
(b) Offline NN-SGF (8) in Algorithm 3.
(c) Online NN-SGF in Algorithm 2.
(d) Neural network trained to emulate $\textsf{U}^{\textsf{lm}}(\boldsymbol{s}_{l},\boldsymbol{\theta})$.
(e) Overvoltage duration.
Figure 4: Highest voltage profile and number of nodes experiencing overvoltages with different optimization methods: (a) batch optimization (BO) using IPOPT; (b) the proposed NN-SGF (8), implemented in a batch configuration as in Figure 1(center); (c) the proposed NN-SGF (8), implemented in an online feedback configuration as in Figure 1(left); (d) method where a neural network is trained to emulate BO solutions (NN-BO). Overvoltage durations for the different strategies are compared in (e); we consider the no control (NC) setup as in Figure 3 and the SGF.
@@
NN-SGF (s2) 60,000 1.7 $\times 10^{-6}$
NN-BO (s3) 80,000 8.3 $\times 10^{-5}$
Figure 4 illustrate how different strategies perform in regulating voltage magnitudes within the bounds $[0.95,1.05]$ p.u.; Figure 4 considers the maximum voltage profile across the system at every time step, as well as the number of nodes that experience overvoltages. These metrics are evaluated over the monitored set $\mathcal{M}=\{93,54,85,84,48,39,91,71,62,67\}$, selected to cover critical voltage locations, such as end-of-feeder buses and buses that approach or violate voltage limits in the baseline no-control case (Fig. 3); we report $\max_{j\in\mathcal{M}}|v_{j}(t)|$ (left axis) together with $N_{\mathrm{imp}}(t):=\bigl|\{\,j\in\mathcal{M}:\ |v_{j}(t)|>\bar{V}\,\}\bigr|$ (secondary axis). The proposed NN–SGF method maintains voltages tightly bounded across all monitored nodes, with only brief and mild excursions slightly above 1.05 p.u. These short-duration deviations are well within the tolerance accepted by distribution utilities and do not compromise protection schemes. The BO method, which solves the full AC OPF offline to convergence, is used as a benchmark. The NN-BO approach, trained to emulate the BO setpoints directly, exhibits significantly more overvoltage excursions than the NN-SGF, reflected in its higher $\max T_{1.05}$ and $\mathrm{mean}\ T_{1.05}$ as shown in 4(e). This confirms that approaches that attempts to learn solutions to the OPF directly cannot ensure feasibility. Importantly, NN-SGF delivers effective online voltage regulation without any iterative optimization and even outperforms widely used schemes such as Volt/Var Control (VVC) and online primal dual methods investigated in [11], which typically suffer from slower response or larger transient violations. Overall, NN-SGF strikes the best balance among voltage compliance, computational efficiency, and system protection, all without introducing operational concerns.
IV-B Computational times
We assess the computational time of the proposed method. In Table II, we first consider an online implementation of the SGF and of the NN-SGF. We recall that the “Online step” refers to the setup in Figure 1(left) where the one evaluation of the SGF (resp., the NN-SGF) is used to generate new setpoints $\boldsymbol{u}(t)$, and the setpoints are sent to the inverters. We averaged the runtime over the full simulation horizon $t\in[\text{06{:}00},\text{20{:}00}]$. Obviously, the computational time of the NN-SGF is much lower, as the SGF involves solving the constrained QP in (11) to obtain the setpoint $\dot{\boldsymbol{u}}(t_{n})=\eta F_{\textsf{in}}(\boldsymbol{u}(t_{n}),\boldsymbol{\xi}(t_{n}),\boldsymbol{\theta}(t_{n}))$.
@@ IV-B Computational times
Overall, an online implementation of the NN-SGF achieves a $\sim\!297\times$ speedup over BO and a $\sim\!45\times$ speedup over the online version of the SGF proposed in [11].
@@
Method Online step Offline implem. Offline
@@
We note that, even when NN–BO is trained using consistent high-voltage local minimizers, the resulting input–output map can be difficult to approximate robustly when deployed in a one-shot, open-loop fashion. The observed performance gap between NN–SGF and NN–BO is therefore not primarily due to set-valuedness in the sense of [37], but rather to how feasibility and approximation errors are handled. In particular, NN–BO applies a learned OPF solution open loop, so that small approximation errors—especially near active constraints—can directly induce voltage violations. By contrast, NN–SGF learns a safe update direction from a strongly convex QP and embeds it in feedback, enforcing (practical) feasibility by construction through forward invariance and explicitly accounting for approximation errors, which leads to improved robustness near constraint boundaries.
We also consider the offline implementation. Here, the “Offline solution” refers to Figure 1(center), where each iteration involves one evaluation of the SGF (resp., the NN-SGF) and one solution to the PF. Again, the PF equations are solved using pandapower, and the average execution time of pandapower was 0.021 seconds. On average, the proposed scheme implemented in an offline fashion required less than 10 iterations, yielding the upper bounds provided in Table II. We note that the proposed online NN-SGF method requires only voltage feedback measurements at buses in the monitored set $\mathcal{M}$. In contrast, BO and NN-BO rely on system-wide measurements or estimates of all non-controllable loads in the network. In practice, acquiring such system-wide information typically requires distribution-level state estimation, which may introduce non-negligible latency beyond the solver runtime. As a result, the end-to-end execution time of BO and NN-BO can be substantially larger in practice than that of NN-SGF [13, 28, 2].
IV-C Cumulative cost
Figure 5 reports the cumulative cost associated with active-power curtailment and reactive-power utilization. Among the considered methods, NN-BO attains the lowest cumulative cost, primarily because it performs less active-power curtailment—thus exporting more available DER generation—and applies smaller reactive-power adjustments, corresponding to weaker corrective actions. However, this apparent advantage comes at the expense of voltage regulation performance. In particular, NN-BO frequently operates outside the admissible voltage region and therefore avoids the curtailment.
In contrast, BO and NN-SGF incur higher cumulative costs because they actively enforce operational constraints. When voltages approach or violate prescribed limits, these methods increase corrective actions through active-power curtailment and/or reactive-power support to drive the system back within allowable bounds. Importantly, NN-SGF enforces feasibility online via feedback-based updates and achieves rapid voltage recovery under fast-changing operating conditions.
Figure 5: Cumulative cost over time for BO, NN–SGF, and NN–BO, where the cost penalizes active-power curtailment and reactive-power utilization.
@@ V Theoretical Analysis
Assumption 2 (Regularity conditions).
Assume that the problem (II-A) is feasible and let $\boldsymbol{u}^{*}$ be an isolated KKT point for (II-A) and a local minimizer, for given $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$. Assume that:
@@
ii) The maps $\boldsymbol{u}\mapsto\phi_{p}(\boldsymbol{u})$, $\boldsymbol{u}\mapsto\phi_{v}(|\boldsymbol{v}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})|)$, $\boldsymbol{u}\mapsto|\boldsymbol{v}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})|$, and $\boldsymbol{u}\mapsto|\boldsymbol{i}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})|$ are twice continuously differentiable over $\mathcal{B}(\boldsymbol{u}^{*},r_{1})$, and their Hessian matrices are positive semi-definite at $\boldsymbol{u}^{*}$.
iii) The Hessian $\nabla^{2}\phi_{p}(\boldsymbol{u}^{*})$ is positive definite. $\Box$
@@
Assumption 3 (Approximation errors).
$\exists~E_{v}<+\infty,E_{J_{v}}<+\infty$ such that $\||\boldsymbol{v}(\boldsymbol{u};\boldsymbol{s}_{l})|-(\boldsymbol{\Gamma}_{v}\boldsymbol{u}+\bar{\boldsymbol{v}}(\boldsymbol{s}_{l}))\|_{\infty}\leq E_{v}$, and $\|\boldsymbol{\Gamma}_{v}-J_{v}(\boldsymbol{u},\boldsymbol{s}_{l})\|\leq E_{J_{v}}$ for any $\boldsymbol{u}\in\mathcal{B}(\boldsymbol{u}^{*},r_{1})$. $\Box$
@@
$\exists~\epsilon_{n}<+\infty$ such that $\|\boldsymbol{n}\|_{\infty}\leq\epsilon_{n}$. Consequently, we let $\epsilon_{v}<+\infty$ and $\epsilon_{i}<+\infty$ such that $\|\boldsymbol{n}_{v}\|_{\infty}\leq\epsilon_{v}$ and $\|\boldsymbol{n}_{i}\|_{\infty}\leq\epsilon_{i}$, respectively. $\Box$
@@
Since the line currents $\boldsymbol{i}$ can be computed from $\boldsymbol{v}$ via Ohm’s Law, Assumption 3 implies that $\exists~E_{i}<+\infty,E_{J_{i}}<+\infty$ such that $\||\boldsymbol{i}(\boldsymbol{u};\boldsymbol{s}_{l})|-(\boldsymbol{\Gamma}_{i}\boldsymbol{u}+\bar{\boldsymbol{i}}(\boldsymbol{s}_{l}))\|_{\infty}\leq E_{i}$ and $\|\boldsymbol{\Gamma}_{i}-J_{i}(\boldsymbol{u},\boldsymbol{s}_{l})\|\leq E_{J_{i}}$ for any $\boldsymbol{u}\in\mathcal{B}(\boldsymbol{u}^{*},r_{1})$. Assumptions 3-4 are motivated by the fact that the error of the linear approximation is small [8, 34], and that in realistic monitoring and SCADA systems, the measurement of the voltage magnitudes are affected by a negligible error [2].
Lastly, Assumption 5 follows from the approximation capabilities of neural networks over compact sets [23, 15].
@@
$:=\arg\min_{\boldsymbol{z}\in\mathbb{R}^{2G}}\|\boldsymbol{z}+\nabla\phi_{p}(\boldsymbol{u})+(\boldsymbol{J}_{v}(\boldsymbol{u};\boldsymbol{s}_{l})+\boldsymbol{\Psi}_{v})^{\top}\nabla\phi_{v}(\boldsymbol{\nu})\|^{2}$
$\hskip 28.45274pt\textrm{s.t.}-(\boldsymbol{J}_{v}(\boldsymbol{u};\boldsymbol{s}_{l})+\boldsymbol{\Psi}_{v})^{\top}\boldsymbol{z}\leq-\beta\left(\mathbf{1}\underline{V}-(|\boldsymbol{v}|+\boldsymbol{n}_{v})\right)$
@@
Assumption 6 (MFCQ and constant-rank conditions).
For any $\boldsymbol{u}\in\mathcal{B}(\boldsymbol{u}^{*},r_{1})$, and any $\boldsymbol{\Psi}_{v}$, $\boldsymbol{\Psi}_{i}$, and $\boldsymbol{n}_{v}$, $\boldsymbol{n}_{i}$ satisfying Assumptions 3-4, the problem (11) is feasible, and satisfies the Mangasarian-Fromovitz Constraint Qualification and the constant-rank condition [26]. $\Box$
@@
Lemma V.1 (Continuity of $F_{\textsf{m}}(\boldsymbol{u},\boldsymbol{n},\boldsymbol{\Psi}_{v},\boldsymbol{\Psi}_{i})$).
Let Assumption 6 hold, and assume that $\boldsymbol{u}\mapsto\phi_{p}(\boldsymbol{u})$, $\boldsymbol{\nu}\mapsto\phi_{v}(\boldsymbol{\nu})$ are twice continuously differentiable over $\mathcal{B}(\boldsymbol{u}^{*},r_{1})$ and for any $\boldsymbol{\nu}$ Then:
@@
Lemma V.1 follows from [26, Theorem 3.6], and by the compactness of the sets $\mathcal{E}_{n}$, $\mathcal{E}_{v}$ and $\mathcal{E}_{i}$. To state the main result, recall that $\boldsymbol{u}^{*}$ is a local optimizer of (II-A). Define $\boldsymbol{\xi}^{*}:=H(\boldsymbol{u}^{*};\boldsymbol{s}_{l})$, $\boldsymbol{J}_{F}:=\frac{\partial F(\boldsymbol{u},H(\boldsymbol{u};\boldsymbol{s}_{l}),\boldsymbol{\theta})}{\partial\boldsymbol{u}}\mid_{\boldsymbol{u}=\boldsymbol{u}^{*}}$, $e_{1}:=-\lambda_{\max}(\boldsymbol{J}_{F})$, and $e_{2}:=-\lambda_{\min}(\boldsymbol{J}_{F})$. Then, from [1], we can write the dynamics as $F(\boldsymbol{u},H(\boldsymbol{u};\boldsymbol{s}_{l}),\boldsymbol{\theta})=\boldsymbol{J}_{F}(\boldsymbol{u}-\boldsymbol{u}^{*})+g(\boldsymbol{u})$, where $g$ satisfies $\|g(\boldsymbol{u})\|\leq L\|\boldsymbol{u}-\boldsymbol{u}^{*}\|^{2}$, $\forall\boldsymbol{u}\in\mathcal{B}(\boldsymbol{u}^{*},r_{2})$, for some $L>0$ and $r_{2}>0$. Define $r:=\min\{r_{1},r_{2}\}$ and $s_{\min}$ as: $s_{\min}=0$ if $r\geq\frac{e_{1}}{L}$, and $s_{\min}=1-\frac{rL}{e_{1}}$ if $r\geq\frac{e_{1}}{L}$. We are ready to state the main result.
Theorem V.2 (Stability).
Consider the OPF problem (II-A) satisfying Assumption 1. Let Assumptions 35 hold for the linear model and the training, and let Assumption 6 hold for (11). Let $\boldsymbol{u}(t)$, $t\geq t_{0}$, be the unique trajectory of (8). Let $\epsilon:=\ell_{J_{v}}E_{J_{v}}+\ell_{J_{i}}E_{J_{i}}+\ell_{n}\epsilon_{n}+\epsilon^{\textsf{NN}}$, and assume that the set $\mathcal{R}:=\left\{s:s_{\min}<s\leq 1,~e_{1}^{-3}e_{2}L\epsilon<s-s^{2}\right\}$ is not empty. Then, for any $s\in\mathcal{R}$, it holds that
@@
Proposition V.3 (Practical feasibility).
@@
$\mathcal{S}_{e,v}$ $:=\{\boldsymbol{u}\in\mathcal{C}:\underline{V}_{e}\leq|v_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{V}_{e},j\in\mathcal{M}\}$ (17a)
$\mathcal{S}_{e,i}$ $:=\{\boldsymbol{u}\in\mathcal{C}:|i_{i}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{I}_{e},j\in\mathcal{L}\}$ (17b)